Denton Company manufactures and sells a single product. Cost data for the product are given:

Variable costs per unit:
Direct materials $ 4
Direct labor 11
Variable manufacturing overhead 3
Variable selling and administrative 1
Total variable cost per unit $ 19
Fixed costs per month:
Fixed manufacturing overhead $ 90,000
Fixed selling and administrative 163,000
Total fixed cost per month $ 253,000

The product sells for $55 per unit. Production and sales data for July and August, the first two months of operations, follow:

Units
Produced
Units
Sold
July 15,000 11,000
August 15,000 19,000

The company’s Accounting Department has prepared the following absorption costing income statements for July and August:

July August
Sales $ 605,000 $ 1,045,000
Cost of goods sold 264,000 456,000
Gross margin 341,000 589,000
Selling and administrative expenses 174,000 182,000
Net operating income $ 167,000 $ 407,000

2. Prepare variable costing income statements for July and August.

3. Reconcile the variable costing and absorption costing net operating incomes.
Answer #1
Direct materials 4
Direct labor 11
Variable manufacturing overhead 3
Variable costing unit product cost 18
2
July August
Sales 605000 1045000
Variable expenses:
Variable cost of goods sold 198000 342000
Variable selling and administrative expenses 11000 19000
Total Variable expenses 209000 361000
Contribution margin 396000 684000
Fixed expenses:
Fixed manufacturing overhead 90000 90000
Fixed selling and administrative expenses 163000 163000
Total Fixed expenses 253000 253000
Net operating income(loss) 143000 431000
3
July August
Variable costing net operating income 143000 431000
Add(deduct) fixed manufacturing overhead deferred in
(released from) inventory under absorption costing
24000 (24000) =4000*6
Absorption costing net operating income 167000 407000
Workings:
Fixed manufacturing overhead per unit 6 =90000/15000
July August
Variable expenses:
Variable cost of goods sold =11000*18 =19000*18
Variable selling and administrative expenses =11000*1 =19000*1